Compute the increase in length of 50m copper wire when its temperature changes from 12 deg C to 32 deg C. For copper, 𝝰 = 1.7 x 10^-5 /deg C.


1
Expert's answer
2021-05-03T10:32:21-0400

The increase in length is


"\Delta L=L_0\alpha \Delta T=(50)(1.7·10^{-5})(32-12)=0.017\text{ m},"

or 1.7 cm.